OpenStudy (anonymous):

Which is an equation of a circle with center (2, -10) and radius 3?

OpenStudy (anonymous):

Equation of a circle: (x-a)^2+(y-b)^2 = r^2 plug and chug.

OpenStudy (anonymous):

a. (x - 2)2 + (y + 10)2 = 3 b.(x - 2)2 + (y + 10)2 = 9 c.(x + 10)2 + (y - 2)2 = 9 d.(x + 2)2 + (y - 10)2 = 3

OpenStudy (anonymous):

thanks